A cyclist is coasting at 11 m/s when she starts down a 450 m long slope that is 30 m high. The cyclist and her bicycle have a combined mass of 70 kg. A steady 14 N drag force due to air resistance acts on her as she coasts all the way to the bottom.
Answer #1

Mass m = 70 kg

u = initial velocity = 11m/s

g = 9.8 m/s^2

F = 14 N

x = 450 m

Apply the Law of conservation of energy and solve the speed.

= 0.5mu^2 + mgH = Fx + 0.5mV^2

Put values..

= 0.5*70*11^2+ 70*9.8*30 = 14*450 +0.5*70*V^2

= 4235+20580=6300+35V^2

= 24815= 6300+35 V^2

=24815-6300= 35V^2

=18515=35V^2

=V^2= 18515/35

=V^2=529

= V=23 m/s